From: NASA HQ/SpaceRef/NASA Watch
Posted: Saturday, September 11, 2004
SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by SpaceRef.com (copyright © 2004) to enhance access to related status reports and NASA activities.
All ISS systems continue to function nominally, except those noted previously or below. Saturday -- and a well-deserved weekend rest day for the crew, which was thanked for yesterday's highly successful PAO events and the Safety Debrief.
Service Module Atmosphere Revitalization Subsystem, Book 2, Mission Operations Directorate, 9 October 2000 [Acrobat] According to this document's introduction "This book contains information for the crew about procedures and rules for the atmosphere revitalization subsystem, Elektron, Vozdukh, Micropurification Unit, and Fire Detection and Suppression Subsystem operations, as well as their schematic and operation logic."
Update on Elektron: The oxygen generator is still off. [Yesterday, CDR Padalka replaced the hydrogen (H2) gas analyzer at the O2 supply line and cleaned residue found when disassembling one of the filters. The unit was restarted, and it operated for about 5 min before shutting down due to bad readings from the newly installed gas analyzer. Another restart was attempted with the gas analyzer inhibited, and Elektron operated but continued to generate the gas analyzer failure message, along with an "H2 Content in Oxygen Higher Limit" message. The crew then shut Elektron down over TsUP/Moscow concerns about the secondary purification unit becoming too warm. For today, procedures for purging the assembly manually with nitrogen (for safety) were uplinked to Padalka, but further troubleshooting has been put off till Monday (9/13) while ground specialists continue their analyses.]
Cabin air was repressurized last night with 10.5 mmHg (Torr) of nitrogen to increase total pressure, which had decreased due to the extensive troubleshooting on Elektron.
After breakfast, the Komandir and his Flight Engineer performed the regular weekly 3-hr. station cleaning. ["Uborka", done every Saturday, includes removal of food waste products, cleaning of compartments with vacuum cleaner, wet cleaning of the Service Module (SM) dining table and other surfaces with "Fungistat" disinfectant and cleaning fan screens to avoid temperature rises.]
Mike Fincke conducted the regular once-a-week maintenance reboot on the operational PCS laptops and the restart of the OCA comm router laptop (every two weeks). [As reported, on 9/9 the OCA (orbital communications adapter) router laptop suffered a failure of its shell that prevented it from booting up. The crew restored the OCA by transferring the router hard drive from the bad shell to the shell of SSC1 (space station computer 1), which has resided in Padalka's "kayuta" in the SM, instead of choosing the SSC6 laptop (MPSD2, Multi-Purpose Support Disc-2) specified by the agreed-upon laptop sparing priority plan. A 30-min task is now being proposed for the crew to move SSC6 into the SM and bring it back online as SSC1, reconfigured for use on the Russian subnet of the Ops LAN and verified for successful connection to the network.]
Gennady Padalka took care of the daily routine maintenance on the SOZh life support system.
Fincke filled out his FFQ (food frequency questionnaire), which keeps an (almost-) regular weekly log of his nutritional intake over time on special MEC (medical equipment computer) software.
Main activity for Mike today was the first part of his chosen "Saturday Science" program, consisting of a session of the MFMG (Miscible Fluids in Microgravity) Thermal Operations Test #2 payload at the Lab's Maintenance Work Area (MWA). The MSG (Microgravity Science Glovebox) is not required. Tomorrow's "Saturday Science" part 2 will focus on software loading on the SNFM (Serial Network Flow Monitor) ER-4 laptop and HRF (Human Research Facility) PC hard drive. [MFMG tests how miscible fluids (fluids that completely dissolve in each other) interact without the interference of gravity using honey and water as the test fluids.]
Working off the Russian discretionary task list, Padalka prepared Matryoshka files stored on the ISS Wiener laptop for downlink to the ground. These files contain logs of the Matryoshka server (BSPN) and are to be used for analyzing BSPN status.
Also from the task list, Gennady was to continued yesterday's brief session with the GFI-8 geophysical Uragan (hurricane) earth-imaging program, using the Kodak 760 DSC (digital still camera) with 800mm-lens from SM windows to take pictures of glaciers of the South Patagonia Ice Field descending into lakes. The raw image files were then to be transferred from the cameras PCMCIA memory card to the TP2 laptop.
Another task-listed job for Gennady was another run of the VC6 "Delta" program's ETD experiment (Investigation of the Coordination of Eye and Head Movements), which he was to conduct before doing any physical exercise. [After a calibration with the calibrating unit, the experiment investigates horizontal eye and head movement coordination, measured Listing's plane, and determined the orientation of the vestibulo-ocular coordinate system, using five target marks on the horizontal plane. Each step required another prior calibration run, using visual target cues or the calibration unit.]
A fourth task list item for Padalka was the removal of the batteries of the two French laptops EGE-1 (at SM window #3) and EGE-2 (at the SM Central Post) due to depletion and loss of capacity. For future ops, the two EGE laptops will only be powered from on-board outlets.
Mike and Gennady completed their daily 2.5-hr. physical exercise program on TVIS, RED exerciser and VELO cycle with bungee cord load trainer.
Weekly Science Update (Expedition Nine -- 20th):
GASMAP: Nothing new.
Human Research Facility/Workstation (HRF WS): Continuing.
Advanced Ultrasound (ADUM): POIC to crew: "Thanks for the great job on Scan Z [on 9/9]. Looking forward to next week’s activities".
Hand Posture Analyzer (HPA: Nothing new.
In-Space Soldering Investigation (ISSE): Nothing new.
Pore Formation and Mobility Investigation (PFMI): Nothing new. Space Acceleration Measurement System (SAMS): Nothing new. Microgravity Acceleration Measurement System (MAMS): Nothing new.
Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES): Nominal.
Protein Crystal Growth Monitoring by Digital Holographic Microscope (PromISS): Nothing new.
Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions (InSPACE): Planned.
Binary Colloidal Alloy Test-3 (BCAT-3): Complete.
Renal Stone (RS): Nothing new.
Synchronized Position Hold, Engage, Reorient, Experimental Satellites (SHERES): Nothing new.
Foot/Ground Reaction Forces during Space Flight (FOOT): Nothing new.
Materials ISS Experiment (MISSE): In progress. Deployed outside on the U.S. Airlock. Nominal and collecting data.
Cellular Biotechnology Support Systems-Fluid Dynamics Investigation (CBOSS-FDI): Nothing new.
Biotechnology Specimen Temperature Controller (BSTC): In planning.
Yeast Group Activation Packs (Yeast GAP): Nothing new.
Earth Knowledge Acquired by Middle School Students (EarthKAM): Successful camera battery charge and function test on 8/28. Thanks!
Earth Science Toward Exploration Research (ESTER): Nothing new.
Serial Network Flow Monitor (SNFM): SNFM completed a 180-min. packet capture on 9/7 (Tuesday) and included a 5-min. HRF (Human Research Facility) downlink. Capture files showed a LAN-1 peak utilization of 14.3% (1,435.7 kbps) and an average of 2.7% (270.4 kbps). HRF utilization was 10.4%. The ground was looking forward to a successful SNFM software load on EXPRESS Rack 4's laptop (ELC-4) tomorrow (9/12).
Fluid Merging Viscosity Measurement (FMVM): Nothing new.
Viscous Liquid Foam--Bulk Metallic Glass (Foam): Nothing new.
BIOPSY (Effects of Prolonged Space Flight on Human Skeletal Muscle): Nothing new.
Coarsening in Solid-Liquid Mixtures-2 (CSLM-2): Planned.
Commercial Generic Bioprocessing Apparatus (CGBA): Nothing new.
Miscible Fluids in Microgravity (MFMG): Nothing new.
Educational Payload Operations (EPO): EPO looks forward to future operations. The crew's previous EPO demonstrations have been well received by NASA Education and payload sponsors.
Capillary Flow Experiment (CFE): Nothing new.
Crew Earth Observations (CEO): Past week’s images of Hurricanes Frances and Ivan in the western Atlantic, Howard in the eastern Pacific, and the Typhoon Songda in the western Pacific, have kept the ground team busy managing the pictures as they are released to the public on their servers. It doesn’t look like the crew missed a single opportunity to photograph them. The views with sun glint are particularly beautiful. Thanks to the crew for their extra effort. A "mysterious" sun glint view of Lake Maracaibo, Venezuela, is being published on the Earth Observatory website this week. Besides the usual sediment and fresh water plumes, oil slicks, and ship wakes on this large, brackish lake, the image depicts the massive, troublesome swirls of Duckweed reported there. It is a truly extraordinary and unexpected shot. ISS' daylight orbit tracks are now swinging “down under” again and targets will be limited for a while. However, with Spring approaching, the lighting and the weather should be much improved from last time.
Today's CEO (crew earth observations) photo targets, limited in the current XPOP attitude by flight rule constraints on the use of the science window, which is available for only ~1/4 of each orbit when not facing forward (in ram), was Biomass burning, SW Brazil (Dynamic event. Fires are burning along most of the perimeter of Brazil’s rainforests and along major routes of human penetration. The result is a vast smoke pall, which appeared both left and right of track. The most meaningful images are those that include smoke free zones to provide contrast with the pall (i.e. margins of the pall, or bounding mountains).
CEO images can be viewed at these websites:
See also the website "Space Station Challenge" at:
To view the latest photos taken by the expedition 9 crew visit:
Major upcoming events:
Expedition 9 Flight Crew Plans can be found at http://spaceflight.nasa.gov/station/timelines/
Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.
ISS Orbit (as of this morning, 10:38am EDT [= epoch]):
ISS Altitude History
Apogee height -- Mean Altitude -- Perigee height
For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html. In addition, information on International Space Station sighting opportunities can be found at http://spaceflight.nasa.gov/realdata/sightings/ on NASA's Human Spaceflight website. The current location of the International Space Station can be found at http://science.nasa.gov/temp/StationLoc.html at NASA's Marshall Space Flight Center. Additional satellite tracking resources can be found at http://www.spaceref.com/iss/tracking.html.
// end //